OpenStudy (anonymous):

solve the equation e^x+25e^-x/2=5 a) what is the solution in terms of natural logarithms b)what is the decimal approximation for the solution

zepdrix (zepdrix):

Is the whole left side being divided by 2, or only the second term?

zepdrix (zepdrix):

Or is the 2 in the exponent? It's not very clear, grr...

OpenStudy (anonymous):

the whole left side is being divided by 2

zepdrix (zepdrix):

\[\Large\rm \frac{e^x+25e^{-x}}{2}=5\]Hmm ok ok I think we can work with this :)

zepdrix (zepdrix):

Multiply both sides by 2,\[\Large\rm e^x+25e^{-x}=10\]This is the tricky step I guess, multiply each side by e^x,\[\Large\rm (e^x)^2+25=10e^x\]Understand those steps so far?

OpenStudy (anonymous):

yes, I understand so far

zepdrix (zepdrix):

If we subtract the 10e^x over to the left side, we actually have a quadratic!\[\Large\rm (e^x)^2-10e^x+25=0\]You can make a substitution from here if you like.. Might make it easier to look at.\[\Large\rm u=e^x\]Plugging that in gives us:\[\Large\rm u^2-10u+25=0\]

zepdrix (zepdrix):

Looks like this one will factor nicely, yes?

OpenStudy (anonymous):

yes, it does

zepdrix (zepdrix):

So what do you get for your factors? c: You'll have just a little bit to do after that. Because we've only found factors for `u` at this point.

OpenStudy (anonymous):

I apologize my head is a little frazzled from doing math for a while

zepdrix (zepdrix):

lol :)

zepdrix (zepdrix):

Successsss! That is our goal in teaching and assigning math work, to frazzle. Learning? Pshh that's an afterthought. Mission accomplished, student successfully frazzled. Ok ok, jk but seriously, how bout them factors? 25 is ummm -5 times -5 right? And -10 is -5 + -5? Those work out, yes?

OpenStudy (anonymous):

lol, yes those work

zepdrix (zepdrix):

\[\Large\rm (u-5)(u-5)=0\]Ok so we get one solution, repeated.\[\Large\rm u=5\]Undoing our substitution gives us:\[\Large\rm e^x=5\]From here, how do we solve for x? Hmmmm, whadya think?

zepdrix (zepdrix):

Comeon Tams! Get your head in the game! We have an exponential... maybeeeee something involving logs? *cough*

OpenStudy (anonymous):

so, would it be ln(e^(5)

zepdrix (zepdrix):

No. Taking the log of each side gives us:\[\Large\rm \ln e^x=\ln 5\]Left side simplifies:\[\Large\rm x=\ln 5\]I'm not sure how your e ended in the log with the 5.. hmm

OpenStudy (anonymous):

I'm not sure either, but was just reworking it on my side and just came to the same solution

zepdrix (zepdrix):

Ah ok :) So that solves part A! For part B just go to your handy calculator. Yay team!

OpenStudy (anonymous):

Thank you very much for your time and assistance

zepdrix (zepdrix):

np c: